1-     A p projectile is launched with an initial seed of 200 m/s at an angle of 60 with respect to the horizontal> compute the range R as measured up the incline


1
Expert's answer
2021-02-01T09:27:48-0500


Given "v=200m\/s"

From the figure above, the horizontal distance, "X=Rcos 20=0.94R"

Vertical distance, "Y=Rsin20=0.342R"

Then, "X=vcos60\times t \implies t=\frac{0.94R}{200cos60}=0.0094R"

Y direction

"Y=v_{y0}t-0.5gt^2"

"0.342R=vsin60\times t-0.5t^2"

"0.342R=200sin60\times 0.0094R-0.5\times 9.81\times 0.0094^2"

"R=2967.2 m"